The current price of a 10 year corporate bod carrying a coupon of 8% is $890. Estimate the bond yield to muturity if thr oer value of the bond is $1000 and the bond pays semi annual coupon payment
Answer #1

Face value 1000 price Coupon 890 (0.08*1000 * 1/2) 40 Maturity period = 10*2 = 20 (because of semiannual payment) we use the